A cubic shipping container had a volume of v3 cubic meters. The height of the container was decreased by a whole number of meters and the width was increased by a whole number of meters so that the volume of the container is now v3 + 2v2 – 3v. Find out the following:
a. By how many meters the height was decreased?
b. By how many meters the width was increased?

Answer by ankor@dixie-net.com(22740)

A cubic shipping container had a volume of v^3 cubic meters.
The height of the container was decreased by a whole number of meters and the
width was increased by a whole number of meters so that the volume of the
container is now v^3 + 2v^2 – 3v. Find out the following:
:
v^3 + 2v^2 – 3v = V
Factor out v
v(v^2 + 2v - 3) = V
v(v + 3)(v - 1) = V
:
a. By how many meters the height was decreased?
1 meter decrease
:
b. By how many meters the width was increased?
3 meter increase